A wandering grizzly bear travels 30 m East, then 40 m in a direction 30 North of East. Calculate its resultant.


1
Expert's answer
2021-09-18T14:59:30-0400

Grizzly's displacement East and North, and then find the resultant displacement and direction:


"D_E=30+40\cos30°=64.6\text{ m}.\\
D_N=40\sin30°=20\text{ m}.\\
D=\sqrt{D_E^2+D_N^2}=67.7\text{ m}.\\\space\\
\theta=\arctan\frac{D_N}{D_E}=17°\text{ N of E}."